OpenStudy (yeval76):

FAN AND MEDAL!! The formula for the circumference of a circle is C = 2πr, where r is the radius. Rearrange the formula to solve for r and select the correct option below. r = C over 2 pi r = 2πC r = C − 2π r = 2 pi over C

OpenStudy (anonymous):

I believe it is A.

OpenStudy (anonymous):

C over 2π

OpenStudy (anonymous):

Ah. I was correct.

OpenStudy (yeval76):

So A is the correct answer? I don't get it. I thought that it was the last one...

OpenStudy (anonymous):

Your answer should be- r= over 2 pi

OpenStudy (anonymous):

*c over 2 pi

OpenStudy (anonymous):

Take it in this say we know 10=5*2 right? So 2=10/5 understand?

OpenStudy (anonymous):

Way*

OpenStudy (yeval76):

Ohhh ok thank you! @joyraheb

OpenStudy (anonymous):

You're welcome :)